The sum of the measures of the angles in any triangle is 180°. Paula has a triangular area in her back yard where she intends to create a garden. If the largest angle is five times the smallest and the middle angle is 20° less than twice the smallest, find the measure of each angle.
Answer by nerdybill(7384)

The sum of the measures of the angles in any triangle is 180°. Paula has a triangular area in her back yard where she intends to create a garden. If the largest angle is five times the smallest and the middle angle is 20° less than twice the smallest, find the measure of each angle.
.
Let x = smallest angle
then
5x = largest angle
2x-20 = middle angle
.
x + 5x + 2x-20 = 180
8x-20 = 180
8x = 200
x = 25 degrees (smallest angle)
.
Largest angle:
5x = 5(25) = 125 degrees
.
Middle angle:
2x-20 = 2(25)-20 = 50-20 = 30 degrees